OpenStudy (anonymous):

Someone please help me solve this equation 8= 4 + d - 5 7/8

OpenStudy (4everaddicted2anime):

the answer is d=9 7/8

OpenStudy (anonymous):

Yes, I know that I got it from an example, but I want to know how to get the answer.

OpenStudy (fortytherapper):

You want to get the numbers on one side, and the variables on the other side (of the equal sign)

OpenStudy (texaschic101):

8 = 4 + d - 5 7/8 8 = 4 + d - 47/8 --- multiply everything by 8 to get rid of fraction 64 = 32 + 8d - 47 64 - 32 + 47 = 8d 79 = 8d 79/8 = d 9 7/8 = d

OpenStudy (4everaddicted2anime):

What I did was subtract the 4 from the right side and the left side to get 4=d-5 7/8 then I added 5 7/8 to both sides d=9 7/8

OpenStudy (anonymous):

Thank you